OpenStudy (anonymous):

solve this inquality: 4a - 3 < - 27

OpenStudy (mathlegend):

4a - 3 = -27

OpenStudy (mathlegend):

4a-3 = -27 +3 +3 4a = -24 -- ---- 4 4 a = -6 a < -6

OpenStudy (mathlegend):

Do you understand?

OpenStudy (mathlegend):

@xsgrriverax